OpenStudy (waheguru):

When WIll This Rocket Hit the Ground y = -5t^2+25t+10

OpenStudy (waheguru):

\[y = -5t^2+25t+10 \] = \[h=-(5t+2)(t-5)\]

OpenStudy (waheguru):

I have gotten the factored form but what Do I do from here to find when it lands ?

OpenStudy (waheguru):

the answer is 5 seconds but I got 5.4 beacuse should we subtrace the tw x-intercets?

OpenStudy (waheguru):

The answer says the rocket will land after 5 seconds but shouldnt it he 5.4 because we need to consider both x-intercets

OpenStudy (mertsj):

Pay attention to the negative sign.

OpenStudy (agent0smith):

\[\Large h=-(5t+2)(t-5) = 0\]You don't subtract the two t values. You just solve that for both t values, one of which will be negative, one positive. One is the t value when it hits the ground. Does it make sense for time to be negative?

OpenStudy (agent0smith):

Solving that gives the x-intercepts, well t-intercepts really. When h=0, that means height is zero, thus it hits the ground.

OpenStudy (waheguru):

We did a similar question in class where the path of water from a water spout was -(d-1)(d-5) and in this case to find the distance traveled (not time) we subtraced 5 -1 giving 4 so thats what made me think to subtract here

OpenStudy (waheguru):

h = -(d-1)(d-5)

OpenStudy (agent0smith):

We aren't finding distance travelled, though. We're finding the time when it hits the ground.

OpenStudy (waheguru):

Also what If both the x-interceptsd were positive?

OpenStudy (waheguru):

for time?

OpenStudy (mertsj):

But I can't figure out where you got those factors. Because they do not multiply back. Did you post the problem incorrectly or something?

OpenStudy (waheguru):

h = -d^2+6d-5

OpenStudy (agent0smith):

d would be distance, t is time, two different variables. You wouldn't get two positive - that would mean the object hits the ground twice (doesn't make much sense), or it's not fired at t=0. And @Mertsj is right, your factoring isn't correct.

OpenStudy (waheguru):

So time cannot have two positive x-intercepts?

OpenStudy (waheguru):

you guys are right, my factoring is wrong

OpenStudy (mertsj):